# VIALM: A Survey and Benchmark of Visually Impaired Assistance with Large Models

### 29 Jan 2024 v1

@@ Abstract
Visually Impaired Assistance (VIA) aims to automatically help visually impaired (VI) handle daily activities. The advancement of VIA primarily depends on developments in Computer Vision (CV) and Natural Language Processing (NLP), both of which exhibit cutting-edge paradigms with large models (LMs). Furthermore, LMs have shown exceptional multimodal abilities to tackle challenging physically-grounded tasks such as embodied robots. To investigate the potential and limitations of state-of-the-art (SOTA) LMs’ capabilities in VIA applications, we present an extensive study for the task of VIA with LMs (VIALM). In this task, given an image illustrating the physical environments and a linguistic request from a VI user, VIALM aims to output step-by-step guidance to assist the VI user in fulfilling the request grounded in the environment. The study consists of a survey reviewing recent LM research and benchmark experiments examining selected LMs’ capabilities in VIA. The results indicate that while LMs can augment VIA, their output cannot be well grounded in reality (i.e., 25.7% GPT-4’s responses) and lacks fine-grained guidance (i.e., 32.1% GPT-4’s responses).
@@ 1 Introduction
Viewing this revolution, we raise a crucial question: can LMs transform the landscape of VIA? We consequently propose a task of Visual Impaired Assistance with Language Models (VIALM). The input of VIALM is an image illustrating the physical environment and a language request from a VI user about a task to complete in the environment. The output is a step-by-step environment-grounded guidance to assist the VI user in fulfilling their request. To illustrate, Figure 1 demonstrates a representative input and output of VIALM.
Figure 1: A sample input and output of VIALM. Its input is a pair of a visual image of the environment (the left image) and a user request in language (the grey box). The yellow box shows the output guidance for VI users to complete the request within the environment (the right image). The output should convey spatial information (blue words) and be comprehensible (green words) for VI users.
To the best of our knowledge, our study is the first to investigate LMs in VIA extensively. It consists of a thorough survey of important LMs’ work, including LLMs, VLMs, and embodied agents, all of which hold potential advancements for VIA. Subsequently, we present a fundamental benchmark for empirical studies by formulating the task following the Visual Question Answering (VQA) setup. It comprises 200 visual environment images with paired questions (user requests) and answers (guidance). The images cover two types of environments: supermarket and home. Based on the benchmark, we examined 6 SOTA VLMs, GPT-4, CogVLM (Wang et al., 2023b), Qwen-VL (Bai et al., 2023), LLaVA (Liu et al., 2023b), MiniGPT-v2 (Chen et al., 2023a), and BLIVE (Hu et al., 2024), to study their zero-shot capabilities in VIA.
The benchmark experiments reveal that the performance of current SOTA LMs is yet to be considered satisfying. We identify two main limitations in the LMs: (1) they struggle to generate environment-grounded guidance (25.7% of GPT-4’s responses), and (2) they lack tactile, fine-grained guidance (32.1% of GPT-4’s responses). Our experiments also indicate potential solutions: (1) addressing the former requires effective visual processing, while (2) the latter necessitates the fusion of tactile modality into language generation.
For future developments, we have open-sourced the resources [^2] https://github.com/YiyiyiZhao/VIALM. The contributions of this work are threefold:
@@ 2 Large Models
Over the past two years, LMs have dramatically reshaped the fields of CV, NLP, and multimodal research. This section offers a comprehensive and current survey of LMs, focusing on a collection of 43 notable models. A chronological representation of the LMs is depicted in Figure 2. It is observed that the years 2021 and 2022 marked advancements in LLMs. This progress was subsequently followed by an upsurge in the development of VLMs and embodied agents in 2023.
@@ 2.1 Large Language Models
Language models are neural networks with text input and output, initially designed for NLP tasks. Large Language Models (LLMs) distinguish themselves by their extensive scale, indicated by the number of network parameters. An important feature of LLMs is their emergent abilities (Wei et al., 2022b), which have been leveraged in reasoning and decision-making (Yao et al., 2023). Following the definition of LLMs in Zhao et al. (2023a), we focus on models with more than 10 billion parameters. In addition, this survey categorizes LLMs based on architectures (Brown et al., 2020) and pretraining tasks (Zeng et al., 2023) and investigates instruction-tuned LLMs (Wei et al., 2022a), systematically summarized in Table 1.
Pretraining Task
The most widely adopted LLM pretraining task is language modeling (Radford et al., 2019), which predicts the next tokens in an autoregressive manner. The success of GPT series (Brown et al., 2020) has demonstrated this approach’s effectiveness, especially when the language model scale is significantly increased and the training corpus is extensively expanded. Consequently, a substantial number of LLMs have adopted it, including GLaM (Du et al., 2022), Gopher (Rae et al., 2021), LaMDA (Thoppilan et al., 2022), PaLM (Chowdhery et al., 2023), Chinchilla (Hoffmann et al., 2022), OPT (Zhang et al., 2022), BLOOM (Scao et al., 2022), Galactica (Taylor et al., 2022), LLaMA (Touvron et al., 2023a), RWKV (Peng et al., 2023), LLaMA-2 (Touvron et al., 2023b), Falcon (Almazrouei et al., 2023) and Mixtral (Jiang et al., 2024). Meanwhile, some LLMs have implemented mixed schemes to enhance language understanding. For example, GLM-130B has introduced autoregressive blank infilling (Zeng et al., 2023), and more recently, PaLM-2 has adopted a mixture of objectives (Anil et al., 2023).
Model Architecture
@@
Instruction-tuned LLMs
@@ 2.2 Large Vision-Language Models
2022.04
NFNet
(Brock et al., 2021)
Chinchilla Cross-attention dense layers Pretraining
@@ 2. Multi-task training
Visual Encoder
Table 2 indicates that ViT is the most preferred visual encoder in current VLMs. Specifically, Qwen-VL implements the standard ViT framework, whereas PaLI-X utilizes the highly scaled-up 22B variant. The ViT-L/14 and ViT-g/14 models, both examined in the BLIP-2 study, are particularly popular. LLaVA has adopted ViT-L/14, while InstructBLIP, BLIVA, and MiniGPT-v2 have selected ViT-g/14. The ViT transformer architecture offers substantial benefits in visual encoding, including scalability in both model and data size (Dehghani et al., 2023), and a more effective integration of visual and language embeddings (Fang et al., 2023).
LLM
In the early stage of the development of VLMs, several models, including BLIP-2, InstructBLIP, and BLIVA, chose the instruction-tuned pioneer FlanT5 as their foundational LLM component. Recently, the LLaMA family, known for its open-source accessibility and impressive performance in NLP tasks, has attracted significant attention. The release of LLaMA-2 is expected to continue and possibly foster this trend. Numerous VLMs, including LLaVA, MiniGPT-4, InstructBLIP, MiniGPT-v2, and CogVLM, have been developed using language models from the LLaMA family, such as Vicuna, Alpava, and LLaMA-2-Chat, as shown in Table 2.
Connector
BLIP-2 introduced the Q-Former component, which was subsequently adopted by InstructBLIP and BLIVA. However, the most common design for connectors currently is a linear projection layer or a multi-layer perceptron(MLP). Models like LLaVA, PaLI-X, MiniGPT-v2, and CogVLM have adopted this design. Similarly, Qwen-VL has adopted the linear layer connection approach, but it improves its functionality by adding cross-attention mechanisms. Comparing MLP with Q-Former, the LLaVA-v1.5 study (Liu et al., 2023a) suggests that the Q-Former may have shortcomings in effectively managing the output length of LLMs.
Training/Tuning
The prevalent pre-training methodology for VLMs generally consists of three phases. Initially, the fundamental pretraining phase utilizes extensive image-text pairs from the Internet to equip the model with general capabilities for visual-language tasks. Subsequently, models undergo multi-task fine-tuning, during which the dataset selection is more carefully curated. The final stage is centered on instruction tuning. It utilizes a limited amount of instructional or conversational data to optimize the model for interactive user engagement. During these stages, the training components of the models may differ. In the initial stage, the connector may update its parameters. However, models typically do not update the LLM parameters at this stage due to their substantial magnitude. In contrast, many VLMs update the LLM parameters during the final instruction tuning phase. Models with frozen LLM include MiniGPT-4 and BLIVA. Although CogVLM does not update its own LLM parameters, it incorporates an additional module into the LLM and updates this module’s parameters. Models that update LLM parameters include LLaVA, Qwen-VL, and MiniGPT-v2.
@@ 2.3 Embodied Agents
LLM-based agents, especially embodied agents, have demonstrated rapid and significant emergence across various scenarios (Xi et al., 2023). The specialized embodied components enable these agents to interact with environments during problem-solving by perceiving the surroundings and performing actions. These actions can either be autonomously carried out by the agent itself or with the aid of additional tools. In this study, we focus on embodied agents that perform actions independently, rather than relying on external tools. This survey analyzes the use of LLMs/VLMs, categorizing their tasks, environments, and design strategies for transforming them into autonomous agents proficient in task-solving. A systematic summary can be found in Table 3.
@@
Table 3: Summary of LLM-based embodied agents. Categorized by environment, ➀ symbolizes simulated enrionments and ➁ the physical world. In LM output, ➀ means mid-level plans, while ➁ indicates low-level robotic actions. Regarding code interface, ➀ is code-free (language-only), and ➁ involves programming code. In LM operations, ➀ is for prompting, ➁ for tuning. It is observed that most embodied agents employ LM prompting for mid-level plans in natural language, which are then translated into low-level robotic actions.
LLM
@@
GPT-3.5 and GPT-4, accessible through API services, are the favored LLMs for developing embodied agents. As shown in Table 3, models such as Inner Monologue (Huang et al., 2022b), ProgPrompt (Singh et al., 2023), LLM-Planner (Song et al., 2022), Matchat (Zhao et al., 2023b), Grounded Decoding (Huang et al., 2023), Voyager (Wang et al., 2023a), GITM (Zhu et al., 2023), and ViLA (Hu et al., 2023) utilize GPT models as the cognitive core for agents’ thinking and planning processes. Additionally, ’Code as Policy’ (Liang et al., 2023) employs CodeX, a derivative of the GPT series focusing on programming. Another prominent LLM is PaLM, which has been chosen by agents like SayCan (Ichter et al., 2022), Grounded Decoding, and RT-2 (Brohan et al., 2023).
Environment and Task
@@
There are two types of environments in which embodied agents operate: simulated and real-world environments. Within simulated settings, agents are trained to execute predefined skills. For instance, Voyager and GITM are designed for playing Minecraft. Additionally, some agents, functioning as virtual robots, utilize platforms like VirtualHome (Puig et al., 2018) for cost-effective training prior to potential deployment in real-world scenarios. Furthermore, numerious embodied agents are able to execute in the physical world. These include SayCan, ’Code as Policy’, Inner Monologue, ProgPrompt, Matcha, PaLM-E, Grounded Decoding, RT-2, ViLA, AutoRT (Ahn et al., 2024), and so on.
From LLM to Agent
@@
To transform LLMs into agents, a typical method is a two-tier framework: the LLM initially generates natural language sequences as high-level plans, later translated into executable low-level actions. Agents using this approach include Inner Monologue, LLM-Planner, Matcha, PaLM-E, Grounded Decoding, ViLA and AutoRT. Alternatively, some models have implemented an end-to-end action generation scheme. Such agents include SayCan, ’Code as Policy’, ProgPrompt, Voyager, GITM, and RT-2. Notably, RT-2 treats action as a modality and has introduced the concept of a visual-language-action model. In addition, while many agents leverage the prompt approach in utilizing LLMs, RT-2 fine-tunes the LLMs for improved performance. Moreover, as next-step action generation is inherently logical, some research conveys a preference for using code instead of natural language to facilitate this process. Agents that have implemented programming code for action generation include ProgPrompt, ’Code as Policy’, and Voyager.
@@ 3 Benchmark Evaluation
In order to explore the question ’Can LMs transform the landscape of VIA?’ and to develop a roadmap for VIALM, we conducted a benchmark evaluation. This assessment involves measuring and comparing the zero-shot performance of LMs, aiming to help readers understand the models’ strengths and weaknesses in the following four key aspects. (1) Viability: Do LMs possess the potential and capabilities necessary for VIA? (2) Environment Adaptability: If LMs are effective for VIA, can they maintain their effectiveness in complex environments? (3) Functionality: Can LMs achieve effective performance in real-world practice? (4) Model Selection: How do different designs of LMs affect performance in VIA?
Evaluation Models
Chen et al. (2023b) demonstrated that end-to-end VLMs outperform separate visual and language pipelines in embodied decision-making tasks. We selected six SOTA end-to-end VLMs for our evaluation, namely GPT-4, CogVLM, MiniGPT, Qwen-VL, LLaVA and BLIVA. The specific versions used were GPT-4, CogVLM-17B-Chat-V1.1, MiniGPT-v2 with LLaMA-2-Chat-7B, Qwen-VL-Chat with Qwen-7B, LLaVA-v1.5-7B, and Bliva-vicuna-7B. All models operated with float16 precision, with the exception of GPT-4.

### 10 Feb 2024 v2

@@ Abstract
Visually Impaired Assistance (VIA) aims to automatically help the visually impaired (VI) handle daily activities. The advancement of VIA primarily depends on developments in Computer Vision (CV) and Natural Language Processing (NLP), both of which exhibit cutting-edge paradigms with large models (LMs). Furthermore, LMs have shown exceptional multimodal abilities to tackle challenging physically-grounded tasks such as embodied robots. To investigate the potential and limitations of state-of-the-art (SOTA) LMs’ capabilities in VIA applications, we present an extensive study for the task of VIA with LMs (VIALM). In this task, given an image illustrating the physical environments and a linguistic request from a VI user, VIALM aims to output step-by-step guidance to assist the VI user in fulfilling the request grounded in the environment. The study consists of a survey reviewing recent LM research and benchmark experiments examining selected LMs’ capabilities in VIA. The results indicate that while LMs can potentially benefit VIA, their output cannot be well environment-grounded (i.e., 25.7% GPT-4’s responses) and lacks fine-grained guidance (i.e., 32.1% GPT-4’s responses).
@@ 1 Introduction
Figure 1: A sample input and output of VIALM. Its input is a pair of a visual image of the environment (the left image) and a user request in language (the grey box). The yellow box shows the output guidance for VI users to complete the request within the environment (the right image). The output should convey environment-grounded information (blue words) and fine-grained (green words) guidance, integrating tactile support (the last sentence) for VI users.
Viewing this revolution, we raise a crucial question: can LMs transform the landscape of VIA? We consequently propose a task of Visual Impaired Assistance with Language Models (VIALM). The input of VIALM is an image illustrating the physical environment and a language request from a VI user about a task to complete in the environment. The output is a step-by-step environment-grounded guidance to assist the VI user in fulfilling their request. This guidance need to be easy to follow (fine-grained) for these users. Particularly, it is crucial to involve tactile guidance as VI users heavily rely on tactile cognition to complement vision loss (Ottink et al., 2022). Figure 1 shows a sample input and output of VIALM.
To the best of our knowledge, our study is the first to investigate LMs in VIA extensively. It consists of a thorough survey of important LMs’ work, including LLMs, VLMs, and embodied agents, all of which hold potential benefits for VIA. Subsequently, we present a fundamental benchmark for empirical studies by formulating the task following the Visual Question Answering (VQA) setup. It comprises 200 visual environment images with paired questions (user requests) and answers (VIA guidance). The images cover two types of environments: supermarket and home. Based on the benchmark, we examined 6 SOTA VLMs, GPT-4, CogVLM (Wang et al., 2023b), Qwen-VL (Bai et al., 2023), LLaVA (Liu et al., 2023b), MiniGPT-v2 (Chen et al., 2023a), and BLIVE (Hu et al., 2024), to study their zero-shot capabilities in VIA.
Benchmark experiments identify two main limitations in current SOTA LMs: (1) inferiority to generate environment-grounded guidance (25.7% of GPT-4’s responses), and (2) a lack of fine-grained guidance (32.1% of GPT-4’s responses), with a shortfall in integrating tactile sensation. However, it is also revealed that visually-focused LMs excel in environment understanding, while those with advanced language models are superior in generating easy-to-follow guidance. To overcome these limitations, we propose potential solutions: (1) improve visual capabilities to enhance environmental grounding, and (2) incorporate tactile modalities into language generation to produce more fine-grained guidance. A promising direction for future research involves synergistically enhancing multimodal capabilities to boost overall effectiveness.
For future developments, we have open-sourced the resources. [^2] https://github.com/YiyiyiZhao/VIALM The contributions of this work are threefold:
@@ 2 Large Models
Over the past two years, LMs have dramatically reshaped the fields of CV, NLP, and multimodal research. This section offers a comprehensive and current survey of LMs, focusing on 43 notable examples, chronologically depicted in Figure 2. It is observed that the years 2021 and 2022 marked advancements in LLMs, followed by an upsurge in the development of VLMs and embodied agents in 2023 and 2024.
@@ 2.1 Large Language Models
Language models are neural networks with text input and output, initially designed for NLP tasks. Large Language Models (LLMs) stand out by their extensive scale, as measured by network parameters. A key feature of LLMs is their emergent abilities (Wei et al., 2022b), which have been leveraged in reasoning and decision-making (Yao et al., 2023). We focus on models with more than 10 billion parameters, in line with the definition in Zhao et al. (2023a). This survey categorizes LLMs based on architectures and pretraining tasks and investigates instruction-tuned LLMs, summarized in Table 1.
Pretraining Tasks.
The predominant pretraining task for LLMs is language modeling (Radford et al., 2019), which predicts the next tokens in an autoregressive manner. The success of GPT series (Brown et al., 2020) has demonstrated this approach’s effectiveness, especially when the language model scale is significantly increased and the training corpus is extensively expanded. Consequently, a substantial number of LLMs have adopted it, including GLaM (Du et al., 2022), Gopher (Rae et al., 2021), LaMDA (Thoppilan et al., 2022), PaLM (Chowdhery et al., 2023), Chinchilla (Hoffmann et al., 2022), OPT (Zhang et al., 2022), BLOOM (Scao et al., 2022), Galactica (Taylor et al., 2022), LLaMA (Touvron et al., 2023a), RWKV (Peng et al., 2023), LLaMA-2 (Touvron et al., 2023b), Falcon (Almazrouei et al., 2023), and Mixtral (Jiang et al., 2024). Meanwhile, some LLMs have implemented mixed schemes to enhance language understanding. For example, GLM-130B has introduced autoregressive blank infilling (Zeng et al., 2023), and more recently, PaLM-2 has adopted a mixture of objectives (Anil et al., 2023).
Model Architectures.
@@
Instruction-tuned LLMs.
@@ 2.2 Large Vision-Language Models
2022.04 NFNet Chinchilla Cross-attention dense layers Pretraining
@@ 2. Multi-task training
Visual Components.
Table 2 indicates that ViT is the most preferred visual encoder in current VLMs. Qwen-VL implements the standard ViT framework, whereas PaLI-X utilizes the highly scaled-up 22B variant. The ViT-L/14 and ViT-g/14 models, both examined in the BLIP-2 study, are particularly popular. LLaVA has adopted ViT-L/14, while InstructBLIP, BLIVA, and MiniGPT-v2 have selected ViT-g/14. The ViT transformer architecture offers substantial benefits in visual encoding, including scalability in both model and data size (Dehghani et al., 2023), and a more effective integration of visual and language embeddings (Fang et al., 2023).
LLM Components.
During early VLM development, models such as BLIP-2 and InstructBLIP adopted the instruction-tuned FlanT5 as the LLM component. Recently, the open-source, high-performing LLaMA family has drawn significant attention, with LLaMA-2’s release likely to amplify this trend. Many VLMs, including LLaVA, MiniGPT-v2, and CogVLM, have adopted language models from the LLaMA family, such as Vicuna, Alpaca, and LLaMA-2-Chat.
Cross-Modal Connectors.
BLIP-2 introduced the Q-Former component, which was subsequently adopted by InstructBLIP and BLIVA. However, the most common design for connectors currently is a linear projection layer or a multi-layer perceptron (MLP). Models like LLaVA, PaLI-X, MiniGPT-v2, and CogVLM have adopted this design. Similarly, Qwen-VL has adopted the linear layer connection approach with additional cross-attention mechanisms. Comparing MLP with Q-Former, the LLaVA-v1.5 study (Liu et al., 2023a) suggests that the Q-Former may have shortcomings in effectively managing the output length of LLMs.
Training Methods.
VLM training typically goes through three phases. Initially, the pretraining phase utilizes extensive image-text pairs from the Internet to equip the model with general multimodal capabilities. Then, models undergo multi-task fine-tuning with more carefully curated datasets. The final stage centers on instruction tuning with a limited amount of instructional or conversational data. During these stages, the training components of the models may differ. In the first stage, the connector may update its parameters. However, models typically do not update the LLM parameters at this stage due to their substantial scale. In contrast, many VLMs update the LLM parameters during the final instruction tuning phase. Models with frozen LLM include BLIVA. Although CogVLM does not update its own LLM parameters, it incorporates an additional module into the LLM and updates this module’s parameters. Models that update LLM parameters include LLaVA, Qwen-VL, and MiniGPT-v2.
@@ 2.3 Embodied Agents
LLM-based agents, especially embodied agents, have demonstrated rapid and significant emergence (Xi et al., 2023). The embodied components enable them to interact with environments by perceiving the surroundings and performing actions. We focus on embodied agents that perform actions independently rather than relying on external tools. This survey analyzes embodied agents by the LLM component, tasks, environments, and design strategies, as summarized in Table 3.
@@
Table 3: Summary of LLM-based embodied agents. Categorized by environment, ➀ symbolizes simulated enrionments and ➁ the physical world. In LM output, ➀ means mid-level plans, while ➁ indicates low-level robotic actions. Regarding code interface, ➀ is code-free (language-only), and ➁ involves programming code. In LM operations, ➀ is for prompting, and ➁ for tuning. It is observed that most embodied agents employ LM prompting for mid-level plans in natural language, which are then translated into low-level robotic actions.
LLM Components.
@@
GPT-3.5 and GPT-4, known for advanced cognitive abilities, are the favored LLMs for developing embodied agents. Models such as Inner Monologue (Huang et al., 2022b), ProgPrompt (Singh et al., 2023), LLM-Planner (Song et al., 2022), Matchat (Zhao et al., 2023b), Grounded Decoding (Huang et al., 2023), Voyager (Wang et al., 2023a), GITM (Zhu et al., 2023), and ViLA (Hu et al., 2023) utilize GPT models as the cognitive core for thinking and planning. Additionally, ‘Code as Policy’ (Liang et al., 2023) employs CodeX, a derivative of the GPT series focusing on programming. Another prominent LLM is PaLM, which has been chosen by agents like SayCan (Ichter et al., 2022), Grounded Decoding, and RT-2 (Brohan et al., 2023).
Environments and Tasks.
@@
There are two types of environments in which embodied agents operate: simulated and real-world environments. Within simulated settings, agents are trained to execute predefined skills. For instance, Voyager and GITM are designed for playing Minecraft. Additionally, some agents, functioning as virtual robots, utilize platforms like VirtualHome (Puig et al., 2018) for cost-effective training prior to potential deployment in real-world scenarios. Furthermore, numerious embodied agents are able to execute in the physical world. These include SayCan, ‘Code as Policy’, Inner Monologue, ProgPrompt, Matcha, PaLM-E, Grounded Decoding, RT-2, ViLA, AutoRT (Ahn et al., 2024), and so on.
LLM-based Agents.
@@
To transform LLMs into agents, a typical method is a two-tier framework: the LLM initially generates natural language sequences as high-level plans, later translated into executable low-level actions. Agents using this approach include Inner Monologue, LLM-Planner, Matcha, PaLM-E, Grounded Decoding, ViLA, and AutoRT. Alternatively, some models have implemented an end-to-end action generation scheme. Such agents include SayCan, ‘Code as Policy’, ProgPrompt, Voyager, GITM, and RT-2. Particularly, RT-2 treats action as a modality and has introduced the concept of a visual-language-action model. In addition, while many agents leverage the prompt approach in utilizing LLMs, RT-2 fine-tunes the LLMs for improved performance. Moreover, as next-step action generation is inherently logical, some research conveys a preference for using code instead of natural language to facilitate this process. Agents that have implemented programming code for action generation include ProgPrompt, ‘Code as Policy’, and Voyager.
@@ 3 Benchmark Evaluation
In order to explore the question ‘Can LMs transform the landscape of VIA?’ and to develop a roadmap for VIALM, we conducted a benchmark evaluation. This assessment involves measuring and comparing the zero-shot performance of LMs, aiming to help readers understand the models’ strengths and weaknesses in the following four key aspects. (RQ1) Fundamental Skills: Do LMs possess the fundamental multimodal capabilities necessary for VIA? (RQ2) Practical Skills: Can LMs achieve effective performance in practice with complex physical environments? (RQ3) Model Selection: How do different designs of LMs affect performance in VIA?
Models and Setup.
Chen et al. (2023b) demonstrated that end-to-end VLMs outperform separate visual and language pipelines in embodied tasks. We selected six SOTA end-to-end VLMs for our evaluation: GPT-4, CogVLM, MiniGPT, Qwen-VL, LLaVA and BLIVA. The specific versions were GPT-4, CogVLM-17B-Chat-V1.1, MiniGPT-v2 with LLaMA-2-Chat-7B, Qwen-VL-Chat with Qwen-7B, LLaVA-v1.5-7B, and Bliva-vicuna-7B. All models operated with float16 precision, with the exception of GPT-4. We explicitly constructed the prompts, requesting LMs to act as a VI assistant and provide grounded guidance to follow with tactile clues.